OpenStudy (anonymous):

An equation is shown below: 4(2x - 5) = 4 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

OpenStudy (anonymous):

x=3 2*3=6-5=1 1*4=4

OpenStudy (anonymous):

thata confusing

OpenStudy (anonymous):

4 dude 4

OpenStudy (anonymous):

Parantheses comes first in this equation

OpenStudy (anonymous):

2*3=6 right? so 6-5 which is 1 Then u have 4 left so 4*1=4

OpenStudy (anonymous):

so how many solutions are there?

OpenStudy (anonymous):

Not sure i just found 1

OpenStudy (anonymous):

but i need to know how many sloutions there are!!

OpenStudy (anonymous):

2 im 80%sure

OpenStudy (anonymous):

Yes 2

OpenStudy (triciaal):

read the question before answering

OpenStudy (anonymous):

but dont yell at me i said 80%

OpenStudy (triciaal):

Part A how many solutions ?

OpenStudy (triciaal):

a linear equation has only one solution

OpenStudy (anonymous):

Simplifying 4(2x + -5) = 4 Reorder the terms: 4(-5 + 2x) = 4 (-5 * 4 + 2x * 4) = 4 (-20 + 8x) = 4 Solving -20 + 8x = 4 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '20' to each side of the equation. -20 + 20 + 8x = 4 + 20 Combine like terms: -20 + 20 = 0 0 + 8x = 4 + 20 8x = 4 + 20 Combine like terms: 4 + 20 = 24 8x = 24 Divide each side by '8'. x = 3 Simplifying x = 3

OpenStudy (triciaal):